# SPRMamba: Surgical Phase Recognition for Endoscopic Submucosal Dissection with Mamba

### v1

@@ Abstract
Endoscopic Submucosal Dissection (ESD) is a minimally invasive procedure initially designed for the treatment of early gastric cancer but is now widely used for various gastrointestinal lesions. Computer-assisted Surgery systems have played a crucial role in improving the precision and safety of ESD procedures, however, their effectiveness is limited by the accurate recognition of surgical phases. The intricate nature of ESD, with different lesion characteristics and tissue structures, presents challenges for real-time surgical phase recognition algorithms. Existing surgical phase recognition algorithms struggle to efficiently capture temporal contexts in video-based scenarios, leading to insufficient performance. To address these issues, we propose SPRMamba, a novel Mamba-based framework for ESD surgical phase recognition. SPRMamba leverages the strengths of Mamba for long-term temporal modeling while introducing the Scaled Residual TranMamba block to enhance the capture of fine-grained details, overcoming the limitations of traditional temporal models like Temporal Convolutional Networks and Transformers. Moreover, a Temporal Sample Strategy is introduced to accelerate the processing, which is essential for real-time phase recognition in clinical settings. Extensive testing on the ESD385 dataset and the cholecystectomy Cholec80 dataset demonstrates that SPRMamba surpasses existing state-of-the-art methods and exhibits greater robustness across various surgical phase recognition tasks.
@@ 1 Introduction
Endoscopic Submucosal Dissection (ESD) is a groundbreaking minimally invasive procedure that has transformed the management of early gastrointestinal cancers. Compared to traditional surgery, ESD offers patients reduced trauma, faster recovery times, and lower rates of cancer recurrence [1]. However, the intricacy of lesion characteristics and the delicate nature of the gastrointestinal tract pose a risk of unexpected complications during ESD, making proficiency in this technique limited to a few experienced endoscopists. With the increasing detection of gastrointestinal cancers through routine gastric endoscopy screenings, there is a growing demand for skilled endoscopists capable of performing ESD. Computer-assisted surgery (CAS) systems are an advanced medical technology that can significantly enhance surgical efficiency and reduce the risk of complications [2]. Surgical phase recognition (SPR) is a critical and challenging task within CAS systems, aimed at identifying surgical activities from surgical videos [3]. In ESD, the phases carry distinct risks, necessitating accurate phase recognition for real-time monitoring, surgical process optimization, context-aware decision support, and early anomaly detection [4]. Additionally, automated SPR can significantly improve postoperative reporting and video indexing, providing valuable educational resources for novice endoscopists [5]. Therefore, developing efficient and accurate video-based SPR algorithms is essential to meet the demands of modern surgical practice and education.
Limited inter-phase variance and high intra-phase variance are the most critical challenges for automatically recognizing surgical phases from video. Even though a lot of surgical phase recognition methods have been proposed to address this issue, for example, Jin et al. [6] introduced TMRNet, which includes a repository to store remote information to learn the relationship between the current frame and previous frames. Czenpiel et al. [7] proposed TeCNO, an enhanced Multi-Stage Temporal Convolutional Network (MS-TCN) [8]. Gao et al. [9] devised Trans-SVNet, addressing fine-grained temporal features loss in TCNs with a compact Transformer model. However, ESD surgical phase recognition is still challenged since their surgical phases are extremely complex and temporally imbalanced. Therefore, applying existing surgical phase recognition methods directly to ESD data leads to performance degradation [10, 11, 12]. A more efficient temporal modeling approach is needed to address the insufficient temporal modeling capabilities of traditional temporal models.
@@
Recently, a state-space model called Mamba has demonstrated substantial potential in modeling temporal contexts. Unlike traditional temporal models, the Mamba model shows higher robustness and accuracy in dealing with complex surgical phase transitions while avoiding the problem of the Transformer’s high computational complexity. Making Mamba well suited for modeling the temporal context in ESD videos. To fully utilize Mamba’s temporal modeling ability, we propose a Mamba-based surgical phase recognition framework called SPRMamba. SPRMamba leverages a ResNet-50 backbone for spatial feature extraction from ESD videos, followed by the application of four LSTContext modules to effectively combine short-term and long-term temporal contexts for phase recognition. Central to SPRMamba is the Scale Residual TranMamba (SRTM) module, which combines Mamba’s long-term temporal modeling capabilities with the Transformer’s ability to capture fine-grained details, addressing the complex phase relationships inherent in ESD surgeries. Furthermore, according to [13], allowing the model to operate on the complete input sequence is more beneficial compared to only accessing a subset of the input. Therefore, considering the quadratic complexity of the transformer and the high memory usage of (where L represents the input sequence length), it is inappropriate to directly apply the transformer to uncut surgical videos. To address this issue, we designed two sampling strategies, a Long-range Sample and a Window Sample, to reduce computational complexity and support the online phase. Our main contributions can be summarized as follows:
@@
Introducing Temporal Sample Strategy (TSS) with Window and Long-range sampling to reduce computational burden and support online phase recognition.
@@ 2.1 Video understanding
In the field of video understanding, researchers are dedicated to developing models that can effectively capture spatiotemporal features to extract semantic information from dynamic scenes. Traditional video understanding methods often rely on 3D Convolutional Neural Networks (3D-CNNs) [14], which extend 2D convolutions to simultaneously process spatial and temporal dimensions. However, 3D-CNNs tend to have high computational complexity, limiting their application in long video analysis.
In the field of video understanding, effectively capturing spatiotemporal features is crucial for extracting semantic information from dynamic scenes. Early approaches, such as 3D Convolutional Neural Networks (3D-CNNs) [14], extended 2D convolutions to process spatial and temporal dimensions simultaneously, achieving reasonable performance in action segmentation tasks. However, the high computational complexity of 3D-CNNs makes them impractical for analyzing long videos, particularly in medical applications like ESD surgery, where procedures can last several hours. To address computational limitations, Temporal Convolutional Networks (TCNs) [15] emerged, which focus on capturing temporal dependencies through one-dimensional convolutions. While TCNs are more efficient and capable of handling longer video sequences, they struggle to model complex action variations and capture the fine-grained details needed for surgical phase recognition, where subtle movements and transitions between phases play a critical role. To address this, Temporal Convolutional Networks (TCNs) [15] were proposed, focusing on capturing temporal information through one-dimensional convolutions, thereby modeling long-term dependencies with lower computational costs. However, TCNs face limitations in handling complex action variations and capturing fine-grained features. Transformer-based models have gained attention in recent years for their ability to model long-range dependencies using self-attention mechanisms [16]. While Transformers achieve state-of-the-art performance in video understanding tasks by capturing spatiotemporal relationships across frames, they are constrained by their quadratic computational complexity, which poses challenges when applying them to long-duration videos in real-time scenarios, such as ESD surgeries. Despite these advancements, the complexity of ESD surgery poses unique challenges for current methods in video understanding. Surgical phases often involve intricate and fine-grained transitions that are difficult to capture with traditional models. In addition, effectively modeling short-term and long-term spatiotemporal dependencies in long videos while maintaining controllable computational costs remains a major challenge.
@@ 2.2 Surgical Phase Recognition
Research in the field of Surgical Phase Recognition mainly focuses on automatically detecting and classifying different phases of surgery by analyzing surgical videos. Early surgical phase recognition methods relied heavily on hand-designed features and statistical learning methods [17, 18, 19, 20, 3, 21, 22, 23]. However, these methods are limited by empirical design features and usually rely on predefined dependencies, making it difficult to accurately capture subtle motion features with strongly nonlinear dynamics, and their performance suffers greatly. Twinanda et al. [3] proposed the EndoNet model, a CNN-based approach that automatically extracts features and performs phase classification from laparoscopic surgical videos. This approach opened the way for SPR research using deep learning. Since then, researchers have sought to enhance SPR by incorporating more advanced temporal modeling techniques. For instance, long short-term memory (LSTM) networks have been widely adopted in models like PhaseNet [24], EndoLSTM [25], and SV-RCNet [26], enabling the capture of both spatial and temporal dependencies. These methods, by combining deep residual networks (ResNet) with LSTM modules, have shown promising improvements in recognizing surgical phases by effectively modeling temporal series data. However, these methods are computationally expensive when dealing with long time series and prone to misclassification when dealing with complex phase transitions. To address these limitations, Transformer-based architectures have been introduced into SPR due to their capability of capturing long-range temporal dependencies. For example, Czempiel et al. [27] proposed a Transformer-based surgical phase recognition method, which significantly improves the model’s performance in complex surgical scenarios through the self-attention mechanism. More recently, Graph Neural Networks (GNNs) have been employed in SPR, with studies such as Padoy et al. [28] demonstrating their effectiveness in modeling complex spatiotemporal relationships between surgical tools and anatomical structures. Although existing techniques have made significant progress in SPR, there are still some challenges, such as the real-time nature of the model, and the accurate recognition of short-time phases. Therefore, this paper aims to construct a novel ESD surgical phase recognition model that is capable of maintaining high accuracy while reducing computational complexity.
@@ 2.3 State Space Models
In the field of vision, [34] proposed Visual Mamba (Vim), which combines position encoding and bi-directional scanning to efficiently capture the global context of an image. Pioneered the application of Mamba in vision tasks. Li et al. [35] constructed a generic framework called Video Mamba Suite to develop, validate, and analyze Mamba’s performance in video understanding. Besides, the great potential of Mamba has inspired a series of works [36, 37, 38, 39, 40], which demonstrated Mamba’s has better performance and higher GPU efficiency than Transformer on visual downstream tasks such as semantic segmentation and video understanding. In our work, we integrate Mamba into the surgical phase recognition model to efficiently capture the temporal context in surgical video.
@@ 3.1 Preliminaries
Structured State Space Sequence Models (S4) and Mamba are discrete versions of continuous systems, which include a time scale parameter that converts continuous parameters and into discrete parameters , . A common conversion method is Zero-Order Hold (ZOH), defined as follows:
@@ 3.2 Scale Residual TranMamba
For the surgical phase recognition task, temporal information plays a crucial role in accurate recognition. However, due to the complexity of ESD surgeries, existing surgical phase recognition algorithms based on traditional temporal models are difficult to apply directly to ESD data. We aim to explore a simple and effective structure combining Mamba that simultaneously models both long-term and short-term temporal contexts. A straightforward approach is to combine the short-term temporal context modeled by the transformer with the long-term temporal context modeled by Mamba. Therefore, we propose the Scaled Residual TranMamba module, as shown in Fig. 1a.
Given the input feature , the SRTM module first applies Instance norm[41]. Then, it uses the TranMamba module to capture long-term and short-term temporal context, thereby generating . In addition, to obtain more comprehensive context information, the fusion of and was achieved through scale residual connections. The fused feature is followed by normalization utilizing Instance Norm and then MLP to learn deeper features. The entire process can be described as follows:
@@
3.3 Temporal Sample Strategy
@@
Due to the quadratic complexity of self-attention blocks, it is impractical to apply attention to long sequences of untrimmed videos. This is because the L of the video sequence is very large, so we need to resample it to achieve modeling of long-term and local temporal contexts, as shown in Fig. 1b.
@@
Temporal Window Sample For temporal window sampling (WS), we divide the sequence into non-overlapping windows of size W. Fig. 1b illustrates the case of W=4. Given that different tasks may have different time-dependent ranges, W is task-specific. We used W=64 in practice. The impact of W was evaluated in Section 4. Instead of modeling the temporal context over the entire sequence of length T, we model the temporal context times, where each , , corresponds to each window.
@@
Temporal Long-range Sample For the temporal long-range sample, we sample the input every G to divide G non-overlapping sequences and model the temporal context information of each sequence. In shown in Fig. 1 b, we model the temporal context information of each of the four downsampled sequences. In general, we model the temporal contexts for G where the and are the same, i.e., . The parameter G provides the flexibility to adjust the sparsity based on the available memory budget, e.g., G=1 corresponds to the case where the attention is applied to the entire sequence. In practice, we use G=64 and evaluate the impact of G in Section 4.
@@
LSTContext Block The top of Fig. 2 illustrate the entire LSTContext block. As in previous work [7], we use a 1D dilated convolution with kernel size 3. This is because the dilated convolution increases the receptive field without the need to increase the parameter number by increasing the kernel size. Where the dilation factor for each layer increases by a factor of 2 and the receptive field exponentially expands as the number of layers increases. Therefore, with a few parameters, we achieved a significantly large receptive field in the temporal sequence, which mitigated model overfitting and effectively promoted the accuracy of surgical phase recognition. The dilated convolution is followed by a Gaussian Error Linear Unit (GELU). In the LSTContext block, we first use the window sample and then the long-range sample, as shown in Fig. 1. Finally, we use a linear layer with residual connectivity to output the features for each frame,
3.4 Overview
The architecture of the proposed SPRMamba is shown in Fig. 2. For a video of length L, we first extract the spatial frame-level feature sequence from a fixed ResNet-50[44], where D=2048 is the spatial dimension, and then SPHMamba uses a linear layer to reduce the feature dimension to 64. As in previous works, we repeated each LSTContext block N times, where the dilation factor of the dilation convolution was increased in each layer. After the first N layers of the LTContext block, we use an additional linear layer to further reduce the dimensionality D to 32. The dimensionality reduction method reduces the number of parameters from 2.49 million to 1.23 million without degrading the accuracy. We also use an additional Conv1D followed by a softmax layer to generate the frame-level class probabilities . We proceed with three additional stages, each consisting of layers of LSTContext blocks. At the beginning of each stage, we reset the dilation factor of the temporal convolution to 1 and compute the frame-level class probabilities after each stage, resulting in a multi-stage loss. We use cross-entropy loss and the mean squared error smoothing loss introduced by [8] for supervised training.
4 Experiments

### v2

@@ Abstract
Endoscopic Submucosal Dissection (ESD) is a minimally invasive procedure initially developed for early gastric cancer treatment and has expanded to address diverse gastrointestinal lesions. While computer-assisted surgery (CAS) systems enhance ESD precision and safety, their efficacy hinges on accurate real-time surgical phase recognition, a task complicated by ESD’s inherent complexity, including heterogeneous lesion characteristics and dynamic tissue interactions. Existing video-based phase recognition algorithms, constrained by inefficient temporal context modeling, exhibit limited performance in capturing fine-grained phase transitions and long-range dependencies. To overcome these limitations, we propose SPRMamba, a novel framework integrating a Mamba-based architecture with a Scaled Residual TranMamba (SRTM) block to synergize long-term temporal modeling and localized detail extraction. SPRMamba further introduces the Hierarchical Sampling Strategy to optimize computational efficiency, enabling real-time processing critical for clinical deployment. Evaluated on the ESD385 dataset and the cholecystectomy benchmark Cholec80, SPRMamba achieves state-of-the-art performance (87.64% accuracy on ESD385, +1.0% over prior methods), demonstrating robust generalizability across surgical workflows. This advancement bridges the gap between computational efficiency and temporal sensitivity, offering a transformative tool for intraoperative guidance and skill assessment in ESD surgery. The code is accessible at https://github.com/Zxnyyyyy/SPRMamba.
@@ 1 Introduction
Endoscopic Submucosal Dissection (ESD) is a minimally invasive procedure for treating early gastrointestinal cancers, offering advantages such as reduced trauma and faster recovery [1]. However, the technical complexity of ESD, due to intricate lesion characteristics and the delicate gastrointestinal tract, limits its widespread adoption despite increasing demand for skilled endoscopists amid rising early cancer detection. Computer-assisted surgery (CAS) systems are an advanced medical technology that can significantly enhance ESD efficiency and reduce the risk of complications [2]. A critical component of CAS is surgical phase recognition (SPR), which enables real-time monitoring, procedural optimization, and surgical education by analyzing video data [3, 5]. Therefore, developing efficient and accurate video-based SPR algorithms is essential to meet the demands of modern surgical practice and education.
Limited inter-phase variance (subtle visual distinctions between phases see Fig.1(a)) and high intra-phase variance (diverse tool motions within a single phase see Fig.1(b)) are the most critical challenges for automatically recognizing surgical phases from video. Recent studies (e.g., TMRNet [6], TeCNO [7], Trans-SVNet [9]) solved this issue by capturing the long-term frame-wise relation between the current frame and previous frames. However, these methods falter in ESD’s unique environment. Because of TCNs [8] lose fine-grained temporal details due to fixed convolutional kernels, while Transformers incur prohibitive computational costs for long, uncut ESD videos. Moreover, existing frameworks prioritize short-term dependencies, neglecting ESD’s heterogeneous phase durations. These limitations exacerbate misclassifications, particularly during transitions between visually similar phases (e.g., incision vs. ESD, see Fig.1(c)) [10, 11, 12].
@@
Recently, a state-space model called Mamba has demonstrated substantial potential in modeling temporal contexts. Unlike traditional temporal models, the Mamba model shows higher robustness and accuracy in dealing with complex surgical phase transitions while avoiding the problem of the Transformer’s high computational complexity. Making Mamba well suited for modeling the temporal context in ESD videos. To fully utilize Mamba’s temporal modeling ability, we propose SPRMamba, a novel framework that integrates Mamba’s linear-complexity state-space modeling with a lightweight Transformer to capture both long-range dependencies and fine-grained temporal features. Furthermore, according to [13], allowing the model to operate on the complete input sequence is more beneficial compared to only accessing a subset of the input. Therefore, considering the quadratic complexity of the transformer and the high memory usage of (where L represents the input sequence length), it is inappropriate to directly apply the transformer to uncut surgical videos. To address this issue, SPRMamba employs a hierarchical sampling Strategy (Long-range and Window Sampling) to reduce computational overhead, support the online phase, and the LSTContext modules to aggregate multi-scale temporal contexts. Our main contributions can be summarized as follows:
@@
We introduce Hierarchical Sampling Strategy with Window and Long-range sampling to reduce computational burden and support online phase recognition.
@@ 2.1 Video understanding
Effective spatiotemporal modeling is foundational for video understanding, yet existing approaches face critical trade-offs in medical applications like ESD. Early 3D-CNNs [14] jointly modeled spatial and temporal features but incurred prohibitive computational costs () for hour-long ESD videos. To address computational limitations, Temporal Convolutional Networks (TCNs) [15] emerged, which focus on capturing temporal dependencies through one-dimensional convolutions. While TCNs are more efficient and capable of handling longer video sequences, they struggle to model complex action variations and capture the fine-grained details needed for surgical phase recognition. Transformer-based models have gained attention in recent years for their ability to model long-range dependencies using self-attention mechanisms [16]. While Transformers achieve state-of-the-art performance in video understanding tasks by capturing spatiotemporal relationships across frames, but introduce quadratic complexity (), limiting real-time deployment. Despite these advancements, the complexity of ESD surgery poses unique challenges for current methods in video understanding. Surgical phases often involve intricate and fine-grained transitions that are difficult to capture with traditional models. In addition, effectively modeling short-term and long-term spatiotemporal dependencies in long videos while maintaining controllable computational costs remains a major challenge.
@@ 2.2 Surgical Phase Recognition
Research in the field of Surgical Phase Recognition mainly focuses on automatically detecting and classifying different phases of surgery by analyzing surgical videos. Early surgical phase recognition methods relied heavily on hand-designed features and statistical learning methods [17, 18, 19, 20, 3, 21, 22, 23]. However, these methods are limited by empirical design features and usually rely on predefined dependencies, making it difficult to accurately capture subtle motion features with strongly nonlinear dynamics, and their performance suffers greatly. Twinanda et al. [3] proposed the EndoNet model, a CNN-based approach that automatically extracts features and performs phase classification from laparoscopic surgical videos. This approach opened the way for SPR research using deep learning. Since then, researchers have sought to enhance SPR by incorporating more advanced temporal modeling techniques. For instance, long short-term memory (LSTM) networks have been widely adopted in models like PhaseNet [24], EndoLSTM [25], and SV-RCNet [26], enabling the capture of both spatial and temporal dependencies. These methods, by combining deep residual networks (ResNet) with LSTM modules, have shown promising improvements in recognizing surgical phases by effectively modeling temporal series data. However, these methods are computationally expensive when dealing with long time series and prone to misclassification when dealing with complex phase transitions. To address these limitations, Transformer-based architectures have been introduced into SPR due to their capability of capturing long-range temporal dependencies. For example, Czempiel et al. [27] proposed a Transformer-based surgical phase recognition method, which significantly improves the model’s performance in complex surgical scenarios through the self-attention mechanism. More recently, Graph Neural Networks (GNNs) have been employed in SPR, with studies such as Padoy et al. [28] demonstrating their effectiveness in modeling complex spatiotemporal relationships between surgical tools and anatomical structures but added architectural complexity, hindering real-time use. These limitations underscore three unresolved challenges: (1) efficient long-sequence modeling, (2) fine-grained temporal resolution, and (3) real-time inference—critical for ESD’s clinical demands. Therefore, this paper aims to construct a novel ESD surgical phase recognition model that is capable of maintaining high accuracy while reducing computational complexity.
@@ 2.3 State Space Models
In the field of vision, [34] proposed Visual Mamba (Vim), which combines position encoding and bi-directional scanning to efficiently capture the global context of an image. Pioneered the application of Mamba in vision tasks. Li et al. [35] constructed a generic framework called Video Mamba Suite to develop, validate, and analyze Mamba’s performance in video understanding. Besides, the great potential of Mamba has inspired a series of works [36, 37, 38, 39, 40], which demonstrated that Mamba has better performance and higher GPU efficiency than Transformer on visual downstream tasks such as semantic segmentation and video understanding. However, no prior work applies SSMs to surgical phase recognition, particularly in ESD’s challenging environment of low inter-phase variance and extreme temporal imbalance. In our work, we integrate Mamba into the surgical phase recognition model to efficiently capture the temporal context in ESD video.
@@ 3.1 Preliminaries
Structured State Space Sequence Models (S4) and Mamba are discrete versions of continuous systems, which include a time scale parameter that converts continuous parameters and into discrete parameters , . common conversion method is Zero-Order Hold (ZOH), defined as follows:
@@
3.2 Overview
The architecture of the proposed SPRMamba is shown in Fig.2. For a video of length , we first extract the spatial frame-level feature sequence from a fixed ResNet-50 [44], where is the spatial dimension, and then SPHMamba uses a linear layer to reduce the feature dimension to 64. As in previous works, we repeated each LSTContext block times, where the dilation factor of the dilation convolution was increased in each layer. In practice, we use and we evaluate the impact of in Section 4. After the first layers of the LTContext block, we use an additional linear layer to further reduce the dimensionality to 32. The dimensionality reduction method reduces the number of parameters from 2.49 million to 1.23 million without degrading the accuracy. We also use an additional Conv1D followed by a softmax layer to generate the frame-level class probabilities .
We proceed with three additional stages, each consisting of layers of LSTContext blocks. At the beginning of each stage, we reset the dilation factor of the temporal convolution to 1 and compute the frame-level class probabilities after each stage, resulting in a multi-stage loss. We use cross-entropy loss and the mean squared error smoothing loss introduced by [8] for supervised training. Inspired by [yi2021asformer], we apply the cross-attention to the LSTContext blocks in stages 2 to 4, instead of using feature for queries and keys for window and long-term context samples, but instead using predictive . We evaluate the impact of the number of stages in Section 4.
@@ 3.3 Scale Residual TranMamba
For the surgical phase recognition task, temporal information plays a crucial role in accurate recognition. However, due to the complexity of ESD surgeries, existing surgical phase recognition algorithms based on traditional temporal models are difficult to apply directly to ESD data. We aim to explore a simple and effective structure combining Mamba, which simultaneously models both long-term and short-term temporal contexts. A straightforward approach is to combine the short-term temporal context modeled by the transformer with the long-term temporal context modeled by Mamba. Therefore, we propose the Scaled Residual TranMamba module, as shown in Fig.3.
Given the input feature , the SRTM module first applies Instance norm [41]. Then, it uses the TranMamba module to capture long-term and short-term temporal context, thereby generating . In addition, to obtain more comprehensive context information, the fusion of and was achieved through scale residual connections. The fused feature is followed by normalization utilizing Instance Norm, and then an MLP to learn deeper features. The entire process can be described as follows:
@@
3.4 Hierarchical Sampling Strateg
@@
Due to the quadratic complexity of self-attention blocks, it is impractical to apply attention to long sequences of untrimmed videos. This is because the L of the video sequence is very large, so we need to resample it to achieve modeling of long-term and local temporal contexts, as shown in Fig.4.
@@
Temporal Window Sample. For temporal window sampling, we divide the sequence into non-overlapping windows of size . Fig.4 illustrates the case of . Given that different tasks may have different time-dependent ranges, W is task-specific. We used in practice. The impact of W was evaluated in Section 4. Instead of modeling the temporal context over the entire sequence of length , we model the temporal context times, where each , , corresponds to each window.
@@
Temporal Long-range Sample. For the temporal long-range sample, we sample the input every to divide non-overlapping sequences and model the temporal context information of each sequence. In shown in Fig.4 b, we model the temporal context information of each of the four downsampled sequences. In general, we model the temporal contexts for where the and are the same, i.e., . The parameter provides the flexibility to adjust the sparsity based on the available memory budget, e.g., corresponds to the case where the attention is applied to the entire sequence. In practice, we use and evaluate the impact of in Section 4.
@@
LSTContext Block. The top of Fig.2 illustrate the entire LSTContext block. As in previous work [7], we use a 1D dilated convolution with kernel size 3. This is because the dilated convolution increases the receptive field without the need to increase the parameter number by increasing the kernel size. Where the dilation factor for each layer increases by a factor of 2, and the receptive field exponentially expands as the number of layers increases. Therefore, with a few parameters, we achieved a significantly larger receptive field in the temporal sequence, which mitigated model overfitting and effectively promoted the accuracy of surgical phase recognition. The dilated convolution is followed by a Gaussian Error Linear Unit (GELU). In the LSTContext block, we first use the window sample and then the long-range sample, as shown in Fig.4. Finally, we use a linear layer with residual connectivity to output the features for each frame,
4 Experiments and Results Analysis